What is the specific latent heat of a
0.75kg object that transfers 251kJ during a change of state?
1
Expert's answer
2020-09-22T15:36:57-0400

Latent heat is the energy per unit of mass that is absorbed or released during a change of state. Therefore,

"L=\frac{Q}{m}=\frac{251kJ}{0.75kg}=335\frac{kJ}{kg}=335\frac{J}{g}"

Answer: "335\frac{J}{g}"